# Query Performance Prediction using Relevance Judgments Generated by Large Language Models

### v1

@@ 7. Analysis
Judging depth analysis. To answer RQ3, as shown in Section 4.3.2, for predicting IDCG, we devise an approximate strategy to use the items in the top ranks of the ranked list that are predicted as relevant by QPP-GenRE to approximate all the relevant items for a query in the corpus. To investigate the impact of ’s value on the quality of the prediction, we investigate the relationship between the QPP quality of predicting nDCG@10 and the judgment depth to answer the following question: What depth of relevance judgment do we need to consider to get a satisfying performance for predicting nDCG@10? In Figure 2, we plot the correlation coefficients between actual nDCG@10 values and nDCG@10 values predicted by QPP-GenRE for different judging depths in {10, 25, 50, 75, 100, 200, 300, 400, 500, 600, 700, 800, 900, 1000} on TREC-DL 19, 20. We also show exact QPP results with depths at 10, 100 and 1000 in Table 2.
@@
Table 2 reveals that, by judging only 10 items in the ranked list of BM25 and ANCE, we can already outperform all the baselines and achieve state-of-the-art QPP quality on half of the evaluation sets we used, e.g., assessing BM25 on TREC-DL 19 and 21, and assessing ANCE on TREC-DL 19. While judging deeper in the ranked list is essential for predicting recall-oriented measures, satisfactory QPP quality is still attainable with a relatively shallow depth. Moreover, Figure 2 illustrates that judging the top 200 items in a ranked list already reaches the saturation point for assessing BM25, while judging less than 100 top items reaches the saturation point for ANCE. We speculate this is because ANCE has better retrieval quality than BM25, and more relevant items would appear earlier in the ranked list of ANCE than BM25; therefore, a shallower judging depth suffices to approximate all relevant items in the corpus. It emphasizes the need to consider retrieval quality when determining the optimal judgment depth for various rankers.
Fine-tuning vs. few-shot in-context learning. To answer RQ4, we analyze the performance of judging relevance and QPP quality of QPP-GenRE equipped with LLaMA-7B (i) trained with PEFT on human relevance labels, and (ii) few-shot prompted (in-context learning) .444We randomly sample human-labeled demonstration examples from the same set used for fine-tuning LLaMA-7B; each example is a triplet (query, passage, relevant/irrelevant); our experiments show that four examples work best, two with relevant passages and two with irrelevant passages. To evaluate the performance of generating relevance judgments, we compute Cohen’s metric to measure the agreement between relevance judgments made by the TREC assessors (i.e., relevance judgments in qrels) and relevance judgments automatically generated by the fine-tuned or few-shot LLM.555Our preliminary experiments show that prompting LLaMA-7B in a zero-shot way yields a very low agreement with TREC assessors (e.g., Cohen’s value 0.001); we omit the result. Faggioli et al. (2023a) reported the relevance judgment agreement in terms of Cohen’s between TREC assessors and GPT-3.5 (text-davinci-003) on TREC-DL 21. We also consider their Cohen’s value for a comparison. Additionally, to further assess the quality of the generated relevance judgments, we compute the Pearson’s correlation coefficients between BM25’s actual RR@10 values and the predicted RR@10 values for BM25 based on the generated relevance judgments.666We do not report the Pearson’s correlation for GPT-3.5 (text-davinci-003) because the relevance judgments generated by Faggioli et al. (2023a) are not available to us. We report the results in Table 3. LLaMA-7B (fine-tuned) shows “fair” agreement with TREC assessors on all datasets, outperforming its few-shot counterpart regarding generating relevance judgments on each dataset. LLaMA-7B (fine-tuned) even exhibits a higher Cohen’s value than the commercial LLM, GPT-3.5 (text-davinci-003). LLaMA-7B (fine-tuned) surpasses its few-shot counterpart on all datasets in terms of Pearson’s . This reveals that fine-tuning is an effective way to improve LLaMA’s quality in generating relevance judgments, which finally translates to better QPP quality.
Dataset | Method | Cohen’s | P- |
TREC-DL 19 | LLaMA-7B (few-shot) | 0.121 | 0.281 |
LLaMA-7B (fine-tuned) | 0.258 | 0.538 | |
TREC-DL 20 | LLaMA-7B (few-shot) | 0.110 | 0.255 |
LLaMA-7B (fine-tuned) | 0.238 | 0.560 | |
TREC-DL 21 | GPT-3.5 (text-davinci-003) (Faggioli et al., 2023a) | 0.260 | - |
LLaMA-7B (few-shot) | 0.140 | 0.237 | |
LLaMA-7B (fine-tuned) | 0.333 | 0.524 | |
TREC-DL 22 | LLaMA-7B (few-shot) | 0.009 | 0.109 |
LLaMA-7B (fine-tuned) | 0.190 | 0.350 |
Data leakage. It is important to ensure LLaMA (Touvron et al., 2023b, a) is not exposed to the relevance labels of the TREC-DL sets during pre-training. Touvron et al. (2023a) clearly show the various sources used for pertaining LLaMA, such as English CommonCrawl, Wikipedia and Stack Exchange. There is no specific overlap with our evaluation set. Moreover, Table 3 shows the poor performance of non-fine-tuned LLaMA in judging relevance, further implying that LLaMA may not have been exposed to the human relevance labels in our evaluation sets during pre-training.
@@
Error analysis. As QPP-GenRE computes QPP based on generated relevance judgments, we analyze QPP errors from the perspective of relevance judgment generation. Figure 3 shows QPP-GenRE’s QPP errors in predicting the performance of BM25 and ANCE in terms of RR@10 on TREC-DL 19, 20; the error is defined as the distance between the RR@10 values predicted by QPP-GenRE and actual RR@10 values, namely “predicted RR@10 minus actual RR@10.” We find that most RR@10 values predicted by QPP-GenRE tend to be smaller than the actual RR@10 values, indicating that QPP-GenRE performs less effectively in identifying relevant items than irrelevant ones from the top of the ranked list. Table 4 shows the confusion matrices that compare relevance judgments made by TREC assessors (i.e., relevance judgments in qrels) and QPP-GenRE on TREC-DL 19, 20. We find that QPP-GenRE tends to wrongly predict some relevant items as irrelevant (false negatives), which provides a further interpretation of the QPP errors we found above. Therefore, reducing false negatives in generating relevance judgments is a potential way to improve QPP quality by QPP-GenRE. We leave this exploration for future work.
@@
Irrelevant | 1749 | 6206 | 1180 | 8957 |
Computational cost analysis. Table 5 shows the online QPP latency of QPP-GenRE and other BERT-based supervised QPP baselines, on TREC-DL 19, on a single NVIDIA A100 GPU. We compute the inference latency when queries are processed individually. For QPP-GenRE, we consider judging depths at 10, 100, and 200; QPP-GenRE can use batch acceleration for judging items for the same query because each item in a ranked list for a query is independent of each other.777qppBERT-PL first splits a ranked list with 100 items into 25 chunks and then predicts the number of relevant items in each chunk. For a fair comparison, we put 25 chunks into one batch for acceleration. Although QPP-GenRE is more expensive than all baselines when predicting one measure due to the much larger parameter size of LLaMA-7B compared to BERT, QPP-GenRE has lower latency compared to some baselines when predicting multiple IR evaluation measures because multiple measures can be derived from the same set of relevance judgments at no additional cost. E.g., while QPP-GenRE is 56% more expensive than M-QPPF for predicting one measure, it becomes more efficient when predicting 2 or more metrics than M-QPPF. Nevertheless, we acknowledge that QPP-GenRE has higher computational costs than supervised QPP methods when predicting a single measure. Conversely, regression-based QPP baselines (NQA-QPP, BERTQPP and M-QPPF) need to train separate models for different IR evaluation metrics. Although qppBERT-PL is not optimized to learn to output one specific IR evaluation measure, qppBERT-PL does not achieve a promising QPP quality (see Table 1 and 2).
We argue that QPP-GenRE’s latency is still much smaller than some high-performing LLM-based rerankers. E.g., Sun et al. (2023b) shows that a GPT-4-based reranker needs 10 API calls (one call takes 3200ms) to rerank 100 items for a query, resulting in 32000ms in total, which is around 20 times worse than QPP-GenRE’s latency with a judging depth of 100. QPP-GenRE can well fit some knowledge-intensive scenarios where QPP quality is prioritized or users may have a higher tolerance level for latency, such as patent search (Lupu et al., 2013), legal search (Tomlinson et al., 2007). Besides using QPP online, QPP can also be used to analyze a search system’s performance in an offline setting (Faggioli et al., 2023d).
@@
QPP-GenRE () | 1566.25 | 1566.25 | 1566.25 | 1566.25 |
QPP-GenRE () | 2845.43 | 2845.43 | 2845.43 | 2845.43 |
@@ 8. Conclusions & Future Work
This paper has proposed a new QPP framework, QPP-GenRE, which models QPP from the perspective of predicting IR measures based on automatically generated relevance judgments. We have devised an approximation strategy for predicting a recall-oriented IR measure, which only judges a limited number of items in a given ranked list for a query, to avoid the cost of traversing the entire corpus to find all relevant items; the approximation strategy also enables us to study into the impact of various judging depths on QPP quality. We have explored the possibility of using open-source LLMs for generating relevance judgments, to ensure scientific reproducibility and avoid data leakage issues. We have explored training LLaMA with parameter-efficient fine-tuning (PEFT) on human-labeled relevance judgments, to improve the generation quality of relevance judgments. Experiments on TREC-DL 19–22 tracks demonstrate that QPP-GenRE significantly surpasses existing QPP approaches, achieving state-of-the-art QPP quality in assessing lexical and neural rankers for both precision- and recall-oriented IR metrics.
@@
Our work also opens new lines of research around our limitations. First, QPP-GenRE is a QPP framework that can be integrated with various methods for judging relevance. We show the success of QPP-GenRE equipped with LLaMA. Exploring various LLM to find the optimal one for relevance judgment generation is beyond the scope of our work. However, in future, we believe it is interesting to investigate QPP-GenRE’s performance integrated with other open-source LLMs (e.g., Zephyr-7B (Tunstall et al., 2023)) or pointwise LLM-based rerankers that utilize label words (Bommasani et al., 2023; Ma et al., 2023b; Zhuang et al., 2023b) as discussed in the related work. Second, we only show QPP-GenRE’s high effectiveness in predicting the two most primary official metrics (RR@10 and nDCG@10) used at TREC DL 19–22 (Craswell et al., 2022, 2021, 2020, 2019). It is worthwhile to consider other metrics at various cutoffs, such as MAP@100, in future. Third, while QPP-GenRE exhibits a promising QPP quality and can be used in scenarios where QPP quality is prioritized and users have a higher tolerance level for latency, e.g., patent search or post analysis, it is worth improving QPP-GenRE’s efficiency in future to widen its scope of applications. We plan to investigate the following ways: using multiple GPU acceleration because judging each item in a ranked list is independent of each other, distilling knowledge from LLMs to smaller language models (Gu et al., 2023), and compressing LLMs by using lower-bit (e.g., 2-bit) quantization (Chee et al., 2023) or using low-rank factorization (Xu et al., 2023).

### v2

@@ 7.1. Judging depth analysis
To answer RQ3, as detailed in Section 4.3.2, for predicting IDCG, we devise an approximation strategy and use the items in the top ranks of the ranked list that are predicted as relevant by QPP-GenRE to approximate all the relevant items for a query in the corpus. To investigate the impact of the value of on the quality of the prediction, we investigate the relationship between the QPP quality of predicting nDCG@10 and the judgment depth to answer the following question: What depth of relevance judgment do we need to consider to get a satisfactory performance for predicting nDCG@10? In Figure 3, we plot the correlation coefficients between actual nDCG@10 values and nDCG@10 values predicted by QPP-GenRE for different judging depths in {10, 25, 50, 75, 100, 200, 300, 400, 500, 600, 700, 800, 900, 1,000} on TREC-DL 19 and 20. We also show exact QPP results with depths at 10, 100 and 1000 in Table 4.
@@
Table 4 reveals that, by judging only 10 items in the ranked list of BM25 and ANCE, we can already outperform all the baselines and achieve state-of-the-art QPP quality on half of the evaluation sets we used, e.g., assessing BM25 on TREC-DL 19 and 21, and assessing ANCE on TREC-DL 19. While judging deeper in the ranked list is essential for predicting recall-oriented measures, satisfactory QPP quality is still attainable with a relatively shallow depth. Moreover, Figure 3 illustrates that judging the top 200 items in a ranked list already reaches the saturation point for assessing BM25, i.e., there is no significant improvement by judging a higher number of items, while judging less than 100 top items reaches the saturation point for ANCE. We speculate that this is because ANCE has better retrieval quality than BM25, and more relevant items would appear earlier in the ranked list of ANCE than BM25; therefore, a shallower judging depth suffices to approximate all relevant items in the corpus. This emphasizes the need to consider retrieval quality when determining the optimal judgment depth for various rankers.
LLM | TREC-DL 19 | TREC-DL 20 | TREC-DL 21 | TREC-DL 22 | ||||
P- | P- | P- | P- | |||||
GPT-3.5 (text-davinci-003) (Faggioli et al., 2023a) | - | - | - | - | 0.260 | - | - | - |
LLaMA-7B (few-shot) | -0.001 | -0.062 | -0.003 | 0.087 | 0.003 | -0.002 | -0.010 | 0.214 |
Llama-3-8B (few-shot) | 0.018 | 0.042 | 0.027 | 0.087 | 0.021 | 0.180 | -0.035 | 0.087 |
Llama-3-8B-Instruct (few-shot) | 0.315 | 0.510 | 0.227 | 0.372 | 0.238 | 0.462 | 0.049 | 0.388 |
LLaMA-7B (fine-tuned) | 0.258 | 0.715 | 0.238 | 0.627 | 0.333 | 0.547 | 0.038 | 0.388 |
Llama-3-8B (fine-tuned) | 0.381 | 0.544 | 0.342 | 0.681 | 0.347 | 0.612 | 0.082 | 0.568 |
Llama-3-8B-Instruct (fine-tuned) | 0.397 | 0.647 | 0.316 | 0.743 | 0.418 | 0.699 | 0.066 | 0.573 |
7.2. Impact of fine-tuning and the choice of LLMs
To answer RQ4, we analyze the impact of fine-tuning and the choice of LLMs on the quality of generated relevance judgments and QPP. We consider LLaMA-7B (Touvron et al., 2023a), Llama-3-8B (AI@Meta, 2024), and Llama-3-8B-Instruct (AI@Meta, 2024) under two settings: (i) trained with PEFT on human relevance labels (following the same fine-tuning setup as in Section 5.6), and (ii) few-shot prompted (in-context learning) .666We randomly sample human-labeled demonstration examples from the same set used for fine-tuning LLMs; each example is a triplet (query, passage, relevant/irrelevant); our experiments show that two examples work best, one with relevant passages and one with irrelevant passages. We do not report the results for a zero-shot setting because our preliminary experiments show that prompting these LLMs in a zero-shot way yields pretty poor performance.
To evaluate the performance of judging relevance, we compute Cohen’s metric to measure the agreement between relevance judgments made by the TREC assessors (i.e., relevance judgments in the qrels) and relevance judgments automatically generated by a fine-tuned or few-shot LLM, on TREC-DL 19–22. Faggioli et al. (2023a) reported the relevance judgment agreement in terms of Cohen’s between TREC assessors and GPT-3.5 (text-davinci-003) on TREC-DL 21; we also consider their Cohen’s value for comparison. To evaluate QPP quality, we compute the Pearson’s correlation coefficients between BM25’s actual nDCG@10 values and those predicted by QPP-GenRE using relevance judgments generated by an LLM, on TREC-DL 19–22.777We do not report the Pearson’s correlation for GPT-3.5 (text-davinci-003) because the relevance judgments generated by Faggioli et al. (2023a) are not available to us. The judging depth is set to 1000 in a ranked list. We report the results in Table 5.
We have two observations. First, fine-tuning improves the quality of relevance judgment generation and QPP for all LLMs on all datasets. Specifically, all fine-tuned LLMs exhibit improved relevance judgment agreement with the TREC assessors on TREC-DL 19–22. After fine-tuning, LLaMA-7B and Llama-3-8B achieve “fair” agreement with the TREC assessors on TREC-DL 19, 20 and 21,888Note that unlike the qrels files for TREC-DL 19, 20, and 21 which are fully manually annotated, the qrels file for TREC-DL 22 is constructed by first detecting near-duplicate items and manually judging only one representative item from each near-duplicate cluster for a given query (Craswell et al., 2021); this difference may result in variation in Cohen’s values of LLMs across TREC-DL 19, 20, 21, and TREC-DL 22. Llama-3-8B-Instruct (fine-tuned) even achieves “moderate” agreement on TREC-DL 21 (a Cohen’s value of 0.418). All fine-tuned LLMs exhibit a higher Cohen’s value than the commercial LLM, GPT-3.5 (text-davinci-003). All fine-tuned LLMs surpass their corresponding few-shot counterpart on all datasets in terms of Pearson’s . This reveals that fine-tuning is an effective way to improve the quality of LLMs in generating relevance judgments, which finally translates to better QPP quality.
Second, newly released and instruction-tuned LLMs generally perform better. Llama-3-8B shows improved relevance judgment agreement and QPP quality compared to LLaMA-7B in most cases. Llama-3-8B-Instruct further enhances relevance judgment generation and QPP quality over both Llama-3-8B and LLaMA-7B across most cases. Notably, Llama-3-8B-Instruct (few-shot) even performs better than or equally as well as LLaMA-7B (fine-tuned) on TREC-DL 19 and 22. This finding implies that with a more effective LLM QPP-GenRE has the potential to achieve imrpoved QPP performance.
7.3. Integrating QPP-GenRE with an LLM-based re-ranker
To show QPP-GenRE’s compatibility with other types of relevance prediction methods instead of directly asking an LLM to explicitly generate explicit relevance judgments, we adapt a state-of-the-art pointwise LLM-based re-ranker, RankLLaMA (Ma et al., 2023a), into a relevance judgment generator, and then integrate QPP-GenRE with the adapted RankLLaMA. Specifically, we translate a re-ranking score into a relevance judgment by applying a threshold: an item is deemed as “relevant” if its re-ranking score meets or exceeds a given threshold value. We analyze Pearson’s and Kendall’s correlation coefficients between BM25’s actual nDCG@10 values and those predicted by QPP-GenRE integrated with RankLLaMA w.r.t. different threshold values on TREC-DL 19 and 20. We employ RepLLaMA (7B) from Tevatron.999https://github.com/texttron/tevatron/tree/main/examples/rankllama RankLLaMA’s re-ranking scores for BM25 range from -12.93 to 89.90 for TREC-DL 19 and from -14.38 to 8.82 for TREC-DL 20. Thresholds are set at intervals of 0.5. The judging depth is set to 1,000 in a ranked list.
We report the results in Figure 4. We find that RepLLaMA achieves the highest QPP quality on both datasets when the threshold is 1. At this particular threshold, RepLLaMA achieves high Pearson’s values of 0.789 and 0.788 on TREC-DL 19 and 20, respectively. These values exceed those of fine-tuned LLaMA-7B, which achieves Pearson’s values of 0.715 and 0.627 on TREC-DL 19 and 20, respectively, as well as Llama-3-8B-Instruct, which achieves Pearson’s values of 0.647 and 0.743 on TREC-DL 19 and 20, respectively (see Figure 5).101010Note that the comparison is not fair because (i) LLaMA-7B, Llama-3-8B-Instruct and all other supervised QPP methods used in this paper are trained on the development set of MS MARCO V1, while RankLLaMA (Ma et al., 2023a) was trained on the training set of MS MARCO V1, which is much larger. (ii) We employ the official version of MS MARCO V1, while RankLLaMA (Ma et al., 2023a) uses the Tevatron version of MS MARCO V1, where passages are enriched with document titles; Lassance and Clinchant (2023) reveal that incorporating titles leads to enhanced ranking performance. This means that a state-of-the-art pointwise LLM-based re-ranker can be adapted into an effective relevance judgment generator. The high QPP quality achieved by RankLLaMA demonstrates QPP-GenRE’s compatibility with other types of relevance prediction methods besides directly using LLMs as relevance judgment generators (i.e., asking an LLM to explicitly generate explicit relevance judgments).
However, compared to directly regarding an LLM as a relevance judgment generator, adapting an LLM-based re-ranker into a relevance judgment generator requires tuning an appropriate threshold. As demonstrated, re-ranking scores are not normalized and their ranges vary across datasets. Directly using a re-ranker as a relevance judgment generator can cause issues in real-world scenarios. Extra calibration work for re-ranking scores might be necessary.
7.4. QPP-GenRE’s interpretability
@@
As QPP-GenRE computes QPP based on generated relevance judgments, we analyze QPP errors from the perspective of relevance judgment generation. Figure 5 shows the QPP errors of QPP-GenRE integrated with LLaMA-7B in predicting the performance of BM25 and ANCE in terms of RR@10 on TREC-DL 19 and 20; the error is defined as the distance between the RR@10 values predicted by QPP-GenRE and actual RR@10 values, namely “predicted RR@10 minus actual RR@10.” We find that most RR@10 values predicted by QPP-GenRE tend to be smaller than the actual RR@10 values, indicating that QPP-GenRE performs less effectively in identifying relevant items than irrelevant ones in the top of the ranked list. Table 6 shows the confusion matrices that compare relevance judgments made by TREC assessors (i.e., relevance judgments in qrels) and QPP-GenRE integrated with LLaMA-7B on TREC-DL 19 and 20. We find that QPP-GenRE tends to wrongly predict some relevant items as irrelevant (false negatives), which provides a further interpretation of the QPP errors we found above. Therefore, reducing false negatives in generating relevance judgments is a potential way to improve the QPP quality of QPP-GenRE. We leave this exploration for future work.
@@
Irrelevant | 1,749 | 6,206 | 1,180 | 8,957 |
7.5. Computational cost analysis
@@
QPP-GenRE () | 1,566.25 | 1,566.25 | 1,566.25 | 1,566.25 |
QPP-GenRE () | 2,845.43 | 2,845.43 | 2,845.43 | 2,845.43 |
Table 7 shows the online QPP latency of QPP-GenRE integrated with LLaMA-7B and other BERT-based supervised QPP baselines, on TREC-DL 19, on a single NVIDIA A100 GPU. We compute the inference latency when queries are processed individually. For QPP-GenRE, we consider judging depths at 10, 100, and 200; QPP-GenRE can use batch acceleration for judging items for the same query because each item in a ranked list for a query is independent of each other.111111qppBERT-PL first splits a ranked list with 100 items into 25 chunks and then predicts the number of relevant items in each chunk. For a fair comparison, we put 25 chunks into one batch for acceleration. Although QPP-GenRE is more expensive than all baselines when predicting one measure due to the much larger parameter size of LLaMA-7B compared to BERT, QPP-GenRE has lower latency compared to some baselines when predicting multiple IR evaluation measures because multiple measures can be derived from the same set of relevance judgments at no additional cost. E.g., while QPP-GenRE is 56% more expensive than M-QPPF for predicting one measure, it becomes more efficient when predicting 2 or more metrics than M-QPPF. Nevertheless, we acknowledge that QPP-GenRE has higher computational costs than supervised QPP methods when predicting a single measure. Conversely, regression-based QPP baselines (NQA-QPP, BERTQPP and M-QPPF) need to train separate models for different IR evaluation metrics. Although qppBERT-PL is not optimized to learn to output one specific IR evaluation measure, qppBERT-PL does not achieve a promising QPP quality (see Table 2 and 4).
We argue that QPP-GenRE’s latency is still much smaller than some high-performing LLM-based re-rankers. E.g., Sun et al. (2023b) show that a GPT-4-based listwise re-ranker needs 10 API calls (one call takes 3,200ms) to re-rank 100 items for a query, resulting in 32,000ms in total, which is around 20 times worse than QPP-GenRE’s latency with a judging depth of 100. QPP-GenRE can well fit some knowledge-intensive professional search scenarios where QPP quality is prioritized or users may have a higher tolerance level for latency, such as patent search (Lupu et al., 2013), legal search (Tomlinson et al., 2007). Besides using QPP online, QPP can also be used to analyze a search system’s performance in an offline setting (Faggioli et al., 2023d).
@@ 8. Conclusions & Future Work
We have proposed a new QPP framework, QPP-GenRE, which models QPP from the perspective of predicting IR evaluation measures based on automatically generated relevance judgments. We have devised an approximation strategy for predicting an IR evaluation measure considering recall, which only judges a limited number of items in a given ranked list for a query, to avoid the cost of traversing the entire corpus to find all relevant items; the approximation strategy also enables us to study into the impact of various judging depths on QPP quality. We have explored using open-source LLMs for generating relevance judgments, to ensure scientific reproducibility. In addition, we have examinged training open-source LLMs with parameter-efficient fine-tuning (PEFT) on human-labeled relevance judgments, to improve the quality of relevance judgment generation and QPP.
Main findings
Experiments on datasets from the TREC-DL 19–22 tracks demonstrate that QPP-GenRE significantly surpasses existing QPP approaches, achieving state-of-the-art QPP quality in assessing lexical and neural rankers for either a precision-oriented IR metric or an IR metric considering recall. Moreover, we have shown that QPP-GenRE has the potential to conduct QPP more accurately when integrated with a more effective LLM, has a good compatibility with other types of relevance prediction methods (e.g., an LLMs-based re-ranker), and exhibits good interpretability.
Broader implications
QPP-GenRE has the potential to facilitate the practical use of QPP. The limited accuracy and interpretability of current QPP methods make them difficult to use in practical applications (Arabzadeh et al., 2024). However, QPP-GenRE demonstrates significantly improved QPP accuracy and better interpretability, enhancing the reliability of QPP results and potentially facilitating the practical use of QPP. Especially, QPP-GenRE has the potential to benefit some knowledge-intensive professional search scenarios, e.g., legal (Tomlinson et al., 2007) or patent search (Lupu et al., 2013). In such scenarios, accurate QPP is prioritized, interpretable QPP results are needed, and users may have a higher tolerance level for latency. QPP-GenRE also has the potential for practical application in commercial search engines: commercial search engines receive many frequent and repeated queries, and QPP-GenRE can improve QPP efficiency by reusing stored relevance judgments for repeated query-item pairs and only generating relevance judgments for new query-item pairs. Moreover, QPP-GenRE can be used to analyze the ranking quality of a search system in an purely offline setting (Faggioli et al., 2023d), where latency is not necessarily an issue.
Limitations and future work
First, we only consider predicting the ranking quality of widely-used lexical and dense retrievers, and have not investigated QPP-GenRE’s bias towards LLMs-based rankers (Ma et al., 2023a). Given that QPP-GenRE is based on LLM-based relevance predictors, it would be particularly interesting to explore QPP-GenRE’s potential biases when it predicts the ranking quality of LLM-based rankers.
Second, QPP-GenRE is a QPP framework that can be integrated with various relevance prediction approaches. We show the success of QPP-GenRE equipped with LLaMA-7B, Llama-3-8B, Llama-3-8B-Instruct as well as a state-of-the-art pointwise LLM-based re-ranker, RankLLaMA (Ma et al., 2023a). Exploring various LLMs to find the optimal one for relevance prediction is beyond the scope of our work. However, in future, we believe it is valuable to investigate QPP-GenRE’s performance integrated with other open-source LLMs (e.g., Mistral (Jiang et al., 2023)) as relevance judgment generators. It is also interesting to adapt pairwise or listwise LLM-based re-rankers into relevance judgment generators and integrate QPP-GenRE with them.
Third, we only show QPP-GenRE’s high effectiveness in predicting two primary metrics (RR@10 and nDCG@10) used at TREC DL 19–22 (Craswell et al., 2022, 2021, 2020, 2019). It is worthwhile to consider other metrics at various cutoffs in future work, e.g., nDCG@20 and MAP@100.
@@
Fourth, while QPP-GenRE exhibits a promising QPP quality and can be used in scenarios where QPP quality is prioritized and users have a higher tolerance level for latency, e.g., patent search or post analysis, it is worth improving QPP-GenRE’s efficiency in future to widen its scope of applications. We plan to investigate (i) the use of multiple GPUs because judging each item in a ranked list is independent of each other, (ii) distilling knowledge from LLMs to smaller language models (Gu et al., 2023), and (iii) compressing LLMs by using lower-bit (e.g., 2-bit) quantization (Chee et al., 2023) or using low-rank factorization (Xu et al., 2023).